# Annotation-free Automatic Music Transcription with Scalable Synthetic Data and Adversarial Domain Confusion

### v2

@@ 3.1 Synthetic audio
Contrary to synthetic audio, real audio often undergoes some degree of compression during the recording and production process. To accommodate this issue, we apply a limiting process to the synthesized audio. We choose limiting over compression due to its higher computational efficiency and its better capability in managing challenging scenarios. The limiting is applied to the normalized audio with a probability of 80%, with the threshold being set randomly from a uniform distribution in the range of 0.5 to 1.0. After completing this process, normalization is applied again.
@@ 3.2 Transcription architecture
$\underset{E,D}{\min}\>\mathrm{CE}(\bm{i}(\bm{x}_{\mathrm{S}}),D(E(\bm{x}_{\mathrm{S}})))+$ $\lambda\,\mathrm{BCE}(0.5,C(E(\bm{x}_{\mathrm{S}})))$
$+$ $\lambda\,\mathrm{BCE}(0.5,C(E(\bm{x}_{\mathrm{R}}))),$ (2)
where $\mathrm{CE}$ denotes cross entropy loss and $\mathrm{BCE}$ denotes binary cross entropy loss. $\bm{x}_{\mathrm{S}}$ and $\bm{x}_{\mathrm{R}}$ are the input mel spectrograms of synthetic and real data, respectively. $E$, $D$, and $C$ are encoder, decoder, and discriminator, respectively. $\bm{i}(\bm{x}_{\mathrm{S}})$ is the one-hot vector indicating ground truth MIDI tokens of input $\bm{x}_{\mathrm{S}}$, and 00 and $1$ are the ground truth labels of the input class (00 for synthetic and $1$ for real). $\lambda$ balances the transcription loss and the domain confusion loss. The value of $\lambda$ is set to 0.01.
The classification loss Eq. (1) updates the discriminator’s parameters to improve the discrimination performance between synthetic and real data. The first term in Eq. (2), the transcription loss, is associated with transcription performance. The second and third terms in Eq. (2) are the adversarial losses that perform domain confusion to make the output of the discriminator closer to 0.5.

### v3

@@ 3.1 Synthetic audio
We also apply limiting to synthesized audio to simulate the dynamics that arise during the recording and production processes of real audio. With an 80% probability, we set a randomly selected limiting threshold between 0.5 and 1.0 for normalized audio. Normalization is then reapplied after this step.
@@ 3.2 Transcription architecture
$\underset{E,D}{\min}$ $\>\mathrm{CE}(\bm{i}(\bm{x}_{\mathrm{S}}),D(E(\bm{x}_{\mathrm{S}})))$
$+\lambda$ $(\mathrm{BCE}(0.5,C(E(\bm{x}_{\mathrm{S}})))+\mathrm{BCE}(0.5,C(E(\bm{x}_{\mathrm{R}})))),$ (2)
where $\mathrm{CE}$ denotes cross entropy loss and $\mathrm{BCE}$ denotes binary cross entropy loss. $\bm{x}_{\mathrm{S}}$ and $\bm{x}_{\mathrm{R}}$ are the input mel spectrograms of synthetic and real data, respectively. $E$, $D$, and $C$ are encoder, decoder, and discriminator, respectively. $\bm{i}(\bm{x}_{\mathrm{S}})$ is the one-hot vector indicating ground truth MIDI tokens of input $\bm{x}_{\mathrm{S}}$, and 00 and $1$ are the ground truth labels of the input class (00 for synthetic and $1$ for real).
Eq. (1), the classification loss, updates the discriminator’s parameters to improve the discrimination performance between synthetic and real data. The first term in Eq. (2), the transcription loss, updates the encoder and decoder’s parameters to ensure that the decoder output approximates the correct MIDI token data, thereby improving transcription performance. The second term in Eq. (2), the adversarial loss, updates the encoder’s parameters to drive the discriminator output toward 0.5, making it more difficult to distinguish between synthesized and real data. The parameter $\lambda$ In Eq. (2) balances the effects of transcription and domain confusion, and was set to 0.01 through preliminary experiments.